OpenStudy (anonymous):

What is the value of u + v + w when u = –10, v = 4, and w = 8? @confluxepic

OpenStudy (confluxepic):

-10+4+8=-6+8=2 The answer is 2.

OpenStudy (anonymous):

Thanks so much. Sorry if I am annoying you with all the questions. :)